OpenStudy (anonymous):

help! plzz ! can someone help me Write the equation of the line which passes through (–4, 2) and is parallel to y = –x + 6 in slope-intercept form.

OpenStudy (anonymous):

since the line has to be parallel, do you know what slope it must have?

OpenStudy (anonymous):

-1?

OpenStudy (anonymous):

good... so m=-1 is your slope, and the given point is (-4, 2).. put these two knowns into point-slope form then put it into slope-intercept form by solving for y.. point-slope form: \(\large y-y_1=m(x-x_1) \)

OpenStudy (anonymous):

Yes I was doing that but, whe I got to y-2=-1(x+4) I dont know if whe I distribute the -2 I should make it positive or negative like this, y-2=-1+-4 or negative?

OpenStudy (anonymous):

on the right side, distribute the -1 first... what do you get?

OpenStudy (anonymous):

I mean-1x+-4

OpenStudy (anonymous):

yes.. that's correct... but simplified, that right side is -1x-4

OpenStudy (anonymous):

or -x-4

OpenStudy (anonymous):

so the plus sign passes to be negative?

OpenStudy (anonymous):

yes... for example, it's like this 5 + -3 = 5 - 3, right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

And then I add 4 in both sides?

OpenStudy (anonymous):

wich is 0?

OpenStudy (anonymous):

no... you want y by itself on the left...

OpenStudy (anonymous):

get rid of that -2 by adding 2 to both sides...

OpenStudy (anonymous):

oh no I have add 2 in both sides

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

correct....:)

OpenStudy (anonymous):

so I have left y=(-x)-(-2) right?

OpenStudy (anonymous):

no... try again...

OpenStudy (anonymous):

I make everything positive? Am always having trouble when I have to change it from posive or from negative

OpenStudy (anonymous):

\(\large y-2=-x-4 \) \(\large y-2\color{red}{+2}=-x-4\color{red}{+2} \) \(\large y=-x-2 \)

OpenStudy (anonymous):

oh I get it , so that is the equation they are asking me for?

OpenStudy (anonymous):

yes...

OpenStudy (anonymous):

Ok thank you so much! Oh I have another question I you could please help me? Do I have to always use the fomula for this kind of problems?